Consider the linear eigenproblem, 𝐴𝑥=𝜆𝑥, for the matrix

D=[1 1 2


2 1 1


1 1 3 ]


1. Solve for the largest eigenvalue by the direct method using the secant method. Let 𝜆^((0))=5 and 𝜆^((1))=4

2. Solve for the eigenvalues by the QR method


1
Expert's answer
2021-11-09T00:09:37-0500

1.

"f(\lambda_0)=f(5)=\begin{vmatrix}
 1-5 & 1&1 \\
 2 & 1-5&1\\
1&1&3-5
\end{vmatrix}=-4\cdot7+5-2=-25"


"f(\lambda_1)=f(4)=\begin{vmatrix}
 1-4 & 1&1 \\
 2 & 1-4&1\\
1&1&3-4
\end{vmatrix}=-3\cdot2+3+5=2"


slope = "\frac{f(\lambda_1)-f(\lambda_0)}{\lambda_1-\lambda_0}=\frac{2+25}{4-5}=-27"


the largest eigenvalue:


"\lambda_2=\lambda_1-\frac{f(\lambda_1)}{slope}=4+2/27=4.074"


2.

"d_1=\begin{pmatrix}
 1 \\
 2 \\
1
\end{pmatrix}" , "d_2=\begin{pmatrix}
 1 \\
 1\\
1
\end{pmatrix}" , "d_3=\begin{pmatrix}
 2 \\
 1 \\
3
\end{pmatrix}"


"||d_1||=\sqrt{1+2^2+1}=2.449"

"q_1=d_1\/||d_1||"

"q_1^T=\begin{pmatrix}
 0.408 & 0.816&0.408 \\
 
\end{pmatrix}"


"d'_2=d_2-(q_1^Td_2)q_1"


"d'_2=\begin{pmatrix}
 1 \\
 1\\
1
\end{pmatrix}-\begin{pmatrix}
 0.408 & 0.816&0.408 \\
 \end{pmatrix}\begin{pmatrix}
 1 \\
 1\\
1
\end{pmatrix}\begin{pmatrix}
 0.408 \\
 0.816\\
0.408
\end{pmatrix}=\begin{pmatrix}
 0.834 \\
 0.334\\
0.834
\end{pmatrix}"


"||d'_2||=1.226"


"q_2^T=d'_2/||d'_2||=\begin{pmatrix}
 0.680 & 0.272&0.680 \\
 
\end{pmatrix}"


"d'_3=d_3-(q_1^Td_3)q_1-(q_2^Td_3)q_2=\begin{pmatrix}
 2 \\
 1\\
3
\end{pmatrix}-\begin{pmatrix}
 0.408 & 0.816&0.408 \\
 \end{pmatrix}\begin{pmatrix}
 2 \\
 1\\
3
\end{pmatrix}\begin{pmatrix}
 0.408 \\
 0.816\\
0.408
\end{pmatrix}--\begin{pmatrix}
 0.680 & 0.272&0.680 \\
 \end{pmatrix}\begin{pmatrix}
 2 \\
 1\\
3
\end{pmatrix}\begin{pmatrix}
 0.680 \\
 0.272\\
0.680
\end{pmatrix}=\begin{pmatrix}
 2- 0.333-0.924\\
 1-0.666-0.074\\
3-0.499-1.387
\end{pmatrix}=\begin{pmatrix}
 0.743\\
 0.260\\
1.114
\end{pmatrix}"


"||d'_3||=1.364"

"q_3=d'_3\/||d'_3||"

"q_3^T=\begin{pmatrix}
 0.545 & 0.191&0.817 \\
 
\end{pmatrix}"


"Q^{(0)}=\begin{pmatrix}
 q_1 & q_2&q_3 \\
 
\end{pmatrix}=\begin{pmatrix}
 0.408 & 0.680&0.545 \\
 0.816 & 0.272&0.191\\
0.408&0.680&0.817
\end{pmatrix}"


for "R^{(0)}" :

"r_{11}=||d'_1||=2.449,r_{22}=||d'_2||=1.226,r_{33}=||d'_3||=1.364"

"r_{12}=q_1^Td_2=1.632,r_{13}=q_1^Td_3=2.856,r_{23}=q_2^Td_3=3.672"


"R^{(0)}=\begin{pmatrix}
 2.449 & 1.632&2.856 \\
 0 & 1.226&3.672\\
0&0&1.364
\end{pmatrix}"


"D^{(1)}=R^{(0)}Q^{(0)}=\begin{pmatrix}
 2.449 & 1.632&2.856 \\
 0 & 1.226&3.672\\
0&0&1.364
\end{pmatrix}\begin{pmatrix}
 0.408 & 0.680&0.545 \\
 0.816 & 0.272&0.191\\
0.408&0.680&0.817
\end{pmatrix}="


"=\begin{pmatrix}
 3.496 & 4.051&4.000 \\
 2.499 & 2.830&3.234\\
0.557&0.928&1.114
\end{pmatrix}"


The diagonal elements in matrix "D^{(1)}" are the first approximation to the eigenvalues of matrix "D".

So,

"\lambda_1=3.496,\lambda_2=2.830,\lambda_3=1.114"